OpenStudy (anonymous):

Papa's Italian Restaurant has collected data about customer sauce orders. It calculated that P(marinara) = 0.74, P(pesto) = 0.62, and P(marinara or pesto) = 0.90. Determine the P(marinara and pesto).

OpenStudy (anonymous):

0.12 0.46 0.16 0.22

ganeshie8 (ganeshie8):

use below :- p(A or B) = P(A) + P(B) - P(A and B)

OpenStudy (anonymous):

that's the problem I don't know how to use that formula can you help me?

ganeshie8 (ganeshie8):

P(marinara or pesto) = P(marinara) + P(pesto) - P(marinara and pesto)

ganeshie8 (ganeshie8):

read the question and plugin given values above

OpenStudy (anonymous):

so p(.74 or .62)= p(.74)+p(.62)-p(.74 and .62)

ganeshie8 (ganeshie8):

NOPE

ganeshie8 (ganeshie8):

P(marinara or pesto) = P(marinara) + P(pesto) - P(marinara and pesto) plugin given values :- 0.9 = 0.74 + 0.62 - P(marinara and pesto)

ganeshie8 (ganeshie8):

0.9 = 0.74 + 0.62 - P(marinara and pesto) 0.9 = 1.36 - P(marinara and pesto) P(marinara and pesto) = 1.36 - 0.9 = ?

OpenStudy (anonymous):

.46

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

thank you so much

ganeshie8 (ganeshie8):

np :)